OpenStudy (kaitlyn_nicole):

Given the function f(x) = x^2 and k = –3, which of the following represents a vertical shift? f(x) + k kf(x) f(x + k) f(kx)

OpenStudy (jdoe0001):

\(\qquad \qquad \qquad \qquad \textit{function transformations} \\ \quad \\ \begin{array}{rllll} % left side templates f(x)=&{\color{purple}{ A}}({\color{blue}{ B}}x+{\color{red}{ C}})+{\color{green}{ D}} \\ \quad \\ y=&{\color{purple}{ A}}({\color{blue}{ B}}x+{\color{red}{ C}})+{\color{green}{ D}} \\ \quad \\ f(x)=&{\color{purple}{ A}}\sqrt{{\color{blue}{ B}}x+{\color{red}{ C}}}+{\color{green}{ D}} \\ \quad \\ f(x)=&{\color{purple}{ A}}\mathbb{R}^{{\color{blue}{ B}}x+{\color{red}{ C}}}+{\color{green}{ D}} \end{array}\qquad \begin{array}{llll} % right side info \bullet \textit{ stretches or shrinks horizontally by } {\color{purple}{ A}}\cdot {\color{blue}{ B}}\\ \bullet \textit{ horizontal shift by }\frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is negative, to the right}\\ \qquad if\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\textit{ is positive, to the left}\\ \bullet \textit{ vertical shift by }{\color{green}{ D}}\\ \qquad if\ {\color{green}{ D}}\textit{ is negative, downwards}\\ \qquad if\ {\color{green}{ D}}\textit{ is positive, upwards} \end{array}\) which one would that be anyway?

OpenStudy (jdoe0001):

notice when a vertical shift occurs, when D is either positive or negative now, check your choices